Two positive charges q are located along the y axis at y = +a and y = -a, respectively. (a) What is the electric field at a distance x along the x axis? (b) What does this become for x >> a? (c) For x << a?
Answer #1

a)

In triangle BOC

Using Pythagorean theorem

r = sqrt(a2 + x2)

also, Cos = x/r = x/sqrt(a2 + x2)

Electric field by each charge is given as

E = kq/r2

E = k q/(a2 + x2)

Net electric field is given as

Enet = 2 E Cos

Enet = 2 (k q/(a2 + x2)) (x/sqrt(a2 + x2))

Enet = 2 kqx / (a2 + x2)3/2

b)

for x>>a

Enet = 2 kqx / x3

Enet = 2 kq / x2

c)

for x <<a

Enet = 2 kqx / a3